OpenStudy (anonymous):

find a n for the arithmetic sequence a12=8, a14=16

OpenStudy (anonymous):

Arithmetic sequence: \[a_m + n = a _{m+1}\] Therefore:\[a_m + 2n = a_{m+2}\] Or:\[a_{12} + 2n = a_{14}\] And so:\[12 + 2n = 16\]\[n = 2\]